# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

### v2

@@ 6.4 A simple example of probablistic programming
By conditioning the generative model on the observed molecule, we restrict the samples to only those molecules that are similar to the specified molecule. The Metropolis-Hastings algorithm then explores the space of possible molecules and generates samples according to their weights (probabilities) based on the conditioned model.
7 Discussion
@@ 7.1 Limitations
The storage efficiency of the should be further explored. The ADT syntax could result in large file sizes for molecules stored as .hs files, however, the use of list comprehension, and the fact that molfiles and the Molecule data type belong to the same complexity class may mean that the ADT representaion is fact smaller than some mol files, for the same degree of expressiveness. By storing a molecule as function, it is possible to significantly reduce the size of the molecule.
@@ 7.2 Future work
The ADT so far presented has the ability to be improved upon, extended, and utilised in empirical experiments in order to verify and test the representation’s utility across a number of domains in cheminformatics.
@@ 7.2.2 Development of powerful, user-friendly library
To make the most of the ADT, a powerful, user-friendly library of functions should to be created. Functions should include the ability to convert between the ADT and other common representaions, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indicies, and there may be differing atomic coordinates. The permuations of a set of indicies, would however form a group, and the group properties of the ADT can be leveraged for this purpose.
@@ 7.2.4 Experimental validation of the ADT
An additional avenue is using equational reasoning to investiage algebraic approaches in molecular modelling [70].
@@ 8 Conclusion
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, preventing nonsensical molecules from being formed, and going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we implemented reaction networks and explore group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
@@ 9.1 Availability of data and materials
All code used to implement the Algebraic Data Type and the experiments in this paper is available via github, at https://github.com/oliverjgoldstein/chemalgprog

### v3

@@ 3.6.1 Meas & Prob Monads
In the presented code, the Prob monad encapsulates a probability distribution by representing a computation as a function that consumes an infinite, lazily generated Tree of random numbers to yield a result. This mechanism threads random choices through the computation by splitting the tree, thereby providing a deterministic view over randomness. Building on this, the Meas monad is defined as a wrapper around a WriterT transformer applied to Prob, which accumulates a log-weight (using the Product (Log Double) type) alongside the result. This dual nature enables the Meas monad to not only sample outcomes (via the sample function) but also to incorporate scoring (via the score function), which adjusts the unnormalized probability measure. As a result, Meas serves as a framework for building and manipulating probabilistic models that can be employed by inference algorithms such as likelihood weighted importance sampling and Metropolis-Hastings.
@@ 3.7 Limitations
The storage efficiency of the should be further explored. The ADT syntax could result in large file sizes for molecules stored as .hs files, however, the use of list comprehension, and the fact that molfiles and the Molecule data type belong to the same complexity class, may mean that the ADT representation may in fact be smaller than some mol files, for the same degree of expressiveness. By storing a molecule as function or properties as higher-order functions, it may be possible to significantly reduce the size of the molecule.
@@ 3.8 Future work
The ADT so far presented has the ability to be improved upon, extended, and utilised in empirical experiments in order to verify and test the representation’s utility across a number of domains in cheminformatics. It would be useful to have parsers from multiple other representations e.g. SMILES strings to this ADT. Currently only parsers for .sdf files have been written.
@@ 3.8.1 Augmentation of the ADT
3.8.2 Dependent Types
Furthermore, to enforce constraints, and guide programmers in development, dependently-typed functional programming languages, as previously defined, can be used. These languages (e.g. [72]) have many advantages including:
  • Eliminating Physically Impossible Molecules: By enforcing valency constraints at the type level, we can prevent the creation of molecules where atoms exceed their natural bonding capabilities (e.g., a carbon atom with five single bonds).
  • Ensuring Charge Balance: Dependent types can encode rules ensuring that molecules do not have unbalanced charges unless explicitly permitted (e.g., handling ionic compounds separately).
  • Constraining Molecular Topology: We can restrict bond formations to valid molecular configurations, preventing unphysical connectivity (e.g., an atom being bonded to itself).
  • Enforcing Electronic Configuration Validity: By parameterizing orbitals and electron shells with dependent types, we can ensure that electrons are correctly distributed within allowed quantum states.
@@ 3.8.3 Development of powerful, user-friendly library
To make the most of the ADT, a powerful, user-friendly library of functions should to be created. This could be parallelised with high performance Haskell libraries such as Accelerate (https://hackage.haskell.org/package/accelerate). Functions should include the ability to convert between the ADT and other common representations, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indices, and there may be differing atomic coordinates. The permuations of a set of indices, would however form a group, and the group properties of the ADT can be leveraged for this purpose.
@@ 3.8.5 Experimental validation of the ADT
An additional avenue is using equational reasoning to investigate algebraic approaches in molecular modeling [76].
@@ 4 Conclusion
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we explored group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
@@ 5.1 Availability of data and materials
All code used to implement the Algebraic Data Type and the experiments in this article is available via github, at https://github.com/oliverjgoldstein/Mol-Rep-Haskell-Bayes. In line with open source molecular software [77], we have submitted a pull request to https://opensourcemolecularmodeling.github.io/.
@@ 5.2 Competing interests
The authors declare no competing interests.
5.3 Funding
Oliver Goldstein is funded by an EPSRC DTP Scholarship.
Sam March is self-funded.
5.4 Authors’ contributions
5.4.1 Oliver Goldstein
Ideated the project and conceptualized the solution in the form of an Algebraic Data Type, wrote the code base of the library with all experiments and extensions and drafted the initial article
5.4.2 Samuel March
Made substantial revisions to the article’s form and content, contributed to the research, and made misc. contributions to the library
5.5 Acknowledgments
My thanks to my supervisor Jeremy, for putting up with me (easier said than done). We are also thankful for discussions at the Algebra of Programming group at Oxford University.